# Topological holography, quantum criticality, and boundary states

### v3

@@ I Introduction
Classification and characterization of phases of matter has been a central theme in condensed matter physics. Traditionally, phases and phase transitions have been described by Landau symmetry breaking theory. However, recent developments in the field have led to the discovery of novel phases of matter that cannot be described by Landau theory. As a prominent example of phases beyond the Landau paradigm, there has been significant progress in the study of gapped quantum phases, including topological orders with long-rang entanglement[1] and various topological phases distinguished by symmetry, such as symmetry-enriched topological (SET)[2, 3, 4] or symmetry-protected topological (SPT) phases[5], depending on whether there are non-trivial fractionalized excitations.
@@ II.1 Boundary states and partition functions
Therefore, the partition function of the (1+1)$d$ theory can be written as a linear combination of the partition function of the topological order $\mathcal{T}$ on the solid torus with an insertion of of Lagrangian algebra $\mathcal{A}$ around $S^{1}$. The construction can be easily generalized to higher genus surfaces to compute the partition functions.
@@ II.2 Symmetries and operators
Local operators on the right boundary can be organized according to their transformations under the fusion category symmetry. As a result, a local operator can be uniquely attached a bulk anyon label $\alpha$ with $w_{\alpha}>0$. Intuitively, the local operator $\mathcal{O}_{\alpha}$ corresponds to a bulk anyon line $\alpha$ that condenses on the left topological boundary, and stretches across the sandwich, as shown in Fig. 4(a). When $\alpha$ is an non-abelian anyon, there may be multiple condensation channels (splitting idempotents[52]) on the left boundary. Each condensation channel $i$ corresponds to an local operator $\mathcal{O}_{\alpha}^{i}$.
@@
When the symmetry is given by a symmetry group $G$, the topological defect lines on the reference boundary are labeled by group elements, and the fusion satisfies the group multiplications (described by $\text{Vec}_{G}$). The local operators correspond to the bulk anyon lines that can condense on the reference boundary. When we choose the reference boundary to be the charge condensed boundary, the local operators corresponds to the pure charges in the quantum double $D(G)$, labeled by $([1],\pi)$. Here $[1]$ denotes the conjugacy class of the identity element and $\pi$ is an irrep of the centralizer of $[1]$, which equals to the group $G$. The Lagrangian algebra is then
@@
and even under $r$ transformation. The other order parameter has two component and is fully faithful under $S_{3}$. Recall that there are two condensation channels for the $C$ anyon line. Let $\mathcal{O}^{1}$ and $\mathcal{O}^{2}$ be the local operators corresponding to the anyon line $C$ with condensation channel $1$ and $2$ on the reference boundary, respectively (and we fix the condensation channel on the right boundary). Using Eq. (7), Eq. (12), and Eq. (13), we have
@@ II.3 Dualities
Each of this anyonic symmetry corresponds to an invertible codimension-1 twist defect $\mathcal{D}$ in the bulk, such that a bulk anyon is permuted when it passing through the defect. In the sandwich construction, we can insert this twist defect parallel to the sandwich, giving rise to a new (1+1)$d$ phase. If the twist defect satisfies a $\mathbb{Z}_{2}$ fusion rule, each of the twist defect of this kind corresponds to a duality between two different $(1+1)$D gapped phases. More specifically, consider two gapped topological gapped boundary conditions $|\mathcal{L}_{1}\rangle$ and $|\mathcal{L}_{2}\rangle$ on the right boundary. The corresponding partition functions are given by $Z_{1}=\langle\mathcal{A}|\mathcal{L}_{1}\rangle$ and $\mathbb{Z}_{2}=\langle\mathcal{A}|\mathcal{L}_{2}\rangle$, where $\langle\mathcal{A}|$ is the left topological boundary. We say that there is a duality between these two gapped phases, if there exist a twist defect $\mathcal{D}$, such that $\mathcal{D}|\mathcal{L}_{1}\rangle=|\mathcal{L}_{2}\rangle$. Written in terms of the partition functions, this means that $Z_{1}^{\prime}=\langle\mathcal{A}|\mathcal{D}|\mathcal{L}_{1}\rangle=Z_{2}$.
@@
We will encounter examples in $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ gauge theory where there exist a twist defect satisfies a $S_{3}$ fusion rule. This kind of twist defects give rise to a triality for the corresponding (1+1)$d$ systems. For two critical boundaries $\Psi_{1}$ and $\Psi_{2}$ related by the twist defect of this kind: $\mathcal{D}|\Psi_{1}\rangle=|\Psi_{2}\rangle$, there is a relation between the partition functions of these two critical theories:
@@ IV.3.1 Duality to $\mathbb{Z}_{4}$: the effect of changing reference boundary
We can also consider a multi-critical point for the $\mathbb{Z}_{4}$ SSB transition. An example is the 4-state Potts model, whose order-disorder transition is described by the $c=1$ $\mathbb{Z}_{4}$ parafermion CFT. We review the $\mathbb{Z}_{N}$ parafermion CFT in Appendix. B. We can write the partition function by using the character $\chi_{lm}$ of the $\mathbb{Z}_{4}$ parafermion CFT:
@@
where $0\leq l\leq 4$, $-l+2\leq m\leq l$ with $l=m$ mod $2$ and identification $(l,m)\sim(4-l,m\pm 4)$. Notice that the $(l,0)$ primary for $l=1,2,3$ are all $\mathbb{Z}_{4}$-symmetric relevant operators, so with only $\mathbb{Z}_{4}$ symmetry the theory describes a multicritical point. This is also consistent with the 4-state Potts model having a much larger microscopic symmetry (i.e. $S_{4}$) than the $\mathbb{Z}_{4}$ clock model.
Via the correspondence, we have a multi-critical DQCP transition Eq. (101) given by the $\mathbb{Z}_{2}$ orbifold of the $\mathbb{Z}_{4}$ parafermion CFT. It is known that the $\mathbb{Z}_{4}$ parafermion CFT can be represented as $\mathrm{U}(1)_{6}/\mathbb{Z}_{2}$, so the DQCP (multicritical) transition is the $\mathrm{U}(1)_{6}$ CFT. Explicit form of the DQCP partition function is given in Appendix B.
@@ Appendix B $\mathbb{Z}_{N}$ parafermion CFT and the DQCP partition function
$\begin{split}Z_{20}&=Z_{m^{2}}+Z_{em^{2}}+Z_{e^{2}m^{2}}+Z_{e^{3}m^{2}}\\ &=\chi_{0,0}\bar{\chi}_{4,0}+\chi_{1,1}\bar{\chi}_{3,1}+\chi_{1,-1}\bar{\chi}_{3,-1}+\chi_{2,2}\bar{\chi}_{2,2}+\chi_{2,0}\bar{\chi}_{2,0}\\ &\quad+\chi_{3,1}\bar{\chi}_{1,1}+\chi_{3,-1}\bar{\chi}_{1,-1}+\chi_{4,2}\bar{\chi}_{4,-2}+\chi_{4,0}\bar{\chi}_{0,0}+\chi_{4,-2}\bar{\chi}_{4,2},\end{split}$
$\begin{split}Z_{21}&=Z_{m^{2}}+iZ_{em^{2}}-Z_{e^{2}m^{2}}-iZ_{e^{3}m^{2}}\\ &=-\chi_{0,0}\bar{\chi}_{4,0}-i\chi_{1,1}\bar{\chi}_{3,1}+i\chi_{1,-1}\bar{\chi}_{3,-1}+\chi_{2,2}\bar{\chi}_{2,2}-\chi_{2,0}\bar{\chi}_{2,0}\\ &\quad-i\chi_{3,1}\bar{\chi}_{1,1}+i\chi_{3,-1}\bar{\chi}_{1,-1}+\chi_{4,2}\bar{\chi}_{4,-2}-\chi_{4,0}\bar{\chi}_{0,0}+\chi_{4,-2}\bar{\chi}_{4,2},\end{split}$
$\begin{split}Z_{22}&=Z_{m^{2}}-Z_{em^{2}}+Z_{e^{2}m^{2}}-Z_{e^{3}m^{2}}\\ &=\chi_{0,0}\bar{\chi}_{4,0}-\chi_{1,1}\bar{\chi}_{3,1}-\chi_{1,-1}\bar{\chi}_{3,-1}+\chi_{2,2}\bar{\chi}_{2,2}+\chi_{2,0}\bar{\chi}_{2,0}\\ &\quad-\chi_{3,1}\bar{\chi}_{1,1}-\chi_{3,-1}\bar{\chi}_{1,-1}+\chi_{4,2}\bar{\chi}_{4,-2}+\chi_{4,0}\bar{\chi}_{0,0}+\chi_{4,-2}\bar{\chi}_{4,2},\end{split}$
$\begin{split}Z_{23}&=Z_{m^{2}}-iZ_{em^{2}}-Z_{e^{2}m^{2}}+iZ_{e^{3}m^{2}}\\ &=-\chi_{0,0}\bar{\chi}_{4,0}+i\chi_{1,1}\bar{\chi}_{3,1}-i\chi_{1,-1}\bar{\chi}_{3,-1}+\chi_{2,2}\bar{\chi}_{2,2}-\chi_{2,0}\bar{\chi}_{2,0}\\ &\quad+i\chi_{3,1}\bar{\chi}_{1,1}-i\chi_{3,-1}\bar{\chi}_{1,-1}+\chi_{4,2}\bar{\chi}_{4,-2}-\chi_{4,0}\bar{\chi}_{0,0}+\chi_{4,-2}\bar{\chi}_{4,2}.\end{split}$ (217)

### v4

@@ I Introduction
Classification and characterization of phases of matter has been a central theme in condensed matter physics. Traditionally, phases and phase transitions have been described by Landau symmetry breaking theory. However, recent developments in the field have led to the discovery of novel phases of matter that cannot be described by Landau theory. As a prominent example of phases beyond the Landau paradigm, there has been significant progress in the study of gapped quantum phases, including topological orders with long-range entanglement[1] and various topological phases distinguished by symmetry, such as symmetry-enriched topological (SET)[2, 3, 4] or symmetry-protected topological (SPT) phases[5], depending on whether there are non-trivial fractionalized excitations.
@@ II.1 Boundary states and partition functions
Therefore, the partition function of the (1+1)$d$ theory can be written as a linear combination of the partition function of the topological order $\mathcal{T}$ on the solid torus with an insertion of Lagrangian algebra $\mathcal{A}$ around $S^{1}$. The construction can be easily generalized to higher genus surfaces to compute the partition functions.
@@ II.2 Symmetries and operators
Local operators on the right boundary can be organized according to their transformations under the fusion category symmetry. As a result, a local operator can be uniquely attached a bulk anyon label $\alpha$ with $w_{\alpha}>0$. Intuitively, the local operator $\mathcal{O}_{\alpha}$ corresponds to a bulk anyon line $\alpha$ that condenses on the left topological boundary, and stretches across the sandwich, as shown in Fig. 4(a). When $\alpha$ is an non-abelian anyon, there may be multiple condensation channels (splitting idempotents[52]) on the left boundary. Each condensation channel $i$ corresponds to a local operator $\mathcal{O}_{\alpha}^{i}$.
@@
When the symmetry is given by a symmetry group $G$, the topological defect lines on the reference boundary are labeled by group elements, and the fusion satisfies the group multiplications (described by $\text{Vec}_{G}$). The local operators correspond to the bulk anyon lines that can condense on the reference boundary. When we choose the reference boundary to be the charge condensed boundary, the local operators correspond to the pure charges in the quantum double $D(G)$, labeled by $([1],\pi)$. Here $[1]$ denotes the conjugacy class of the identity element and $\pi$ is an irrep of the centralizer of $[1]$, which equals to the group $G$. The Lagrangian algebra is then
@@
and even under $r$ transformation. The other order parameter has two components and is fully faithful under $S_{3}$. Recall that there are two condensation channels for the $C$ anyon line. Let $\mathcal{O}^{1}$ and $\mathcal{O}^{2}$ be the local operators corresponding to the anyon line $C$ with condensation channel $1$ and $2$ on the reference boundary, respectively (and we fix the condensation channel on the right boundary). Using Eq. (7), Eq. (12), and Eq. (13), we have
@@ II.3 Dualities
Each of this anyonic symmetry corresponds to an invertible codimension-1 twist defect $\mathcal{D}$ in the bulk, such that a bulk anyon is permuted when it passes through the defect. In the sandwich construction, we can insert this twist defect parallel to the sandwich, giving rise to a new (1+1)$d$ phase. If the twist defect satisfies a $\mathbb{Z}_{2}$ fusion rule, each twist defect of this kind corresponds to a duality between two different $(1+1)$D gapped phases. More specifically, consider two gapped topological gapped boundary conditions $|\mathcal{L}_{1}\rangle$ and $|\mathcal{L}_{2}\rangle$ on the right boundary. The corresponding partition functions are given by $Z_{1}=\langle\mathcal{A}|\mathcal{L}_{1}\rangle$ and $\mathbb{Z}_{2}=\langle\mathcal{A}|\mathcal{L}_{2}\rangle$, where $\langle\mathcal{A}|$ is the left topological boundary. We say that there is a duality between these two gapped phases, if there exists a twist defect $\mathcal{D}$, such that $\mathcal{D}|\mathcal{L}_{1}\rangle=|\mathcal{L}_{2}\rangle$. Written in terms of the partition functions, this means that $Z_{1}^{\prime}=\langle\mathcal{A}|\mathcal{D}|\mathcal{L}_{1}\rangle=Z_{2}$.
@@
We will encounter examples in $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ gauge theory where there exists a twist defect that satisfies a $S_{3}$ fusion rule. This kind of twist defects give rise to a triality for the corresponding (1+1)$d$ systems. For two critical boundaries $\Psi_{1}$ and $\Psi_{2}$ related by the twist defect of this kind: $\mathcal{D}|\Psi_{1}\rangle=|\Psi_{2}\rangle$, there is a relation between the partition functions of these two critical theories:
@@ IV.3.1 Duality to $\mathbb{Z}_{4}$: the effect of changing reference boundary
We can also consider alternative critical theories for $\mathbb{Z}_{4}$ SSB transitions. An example is $c=1$ $\mathbb{Z}_{4}$ parafermion CFT, which can describe the order-disorder transition in e.g. the Ashkin-Teller model. We review the $\mathbb{Z}_{N}$ parafermion CFT in Appendix. B. We can write the partition function by using the character $\chi_{lm}$ of the $\mathbb{Z}_{4}$ parafermion CFT (using the coset representation ${\rm SU}(2)_{4}/\mathrm{U}(1)$):
@@
where $0<l\leq 4$, $-l+2\leq m\leq l$ with $l=m$ mod $2$ and identification $(l,m)\sim(4-l,m\pm 4)$. There are two $\mathbb{Z}_{4}$ symmetric primary operators: $(2,0)$, which is a relevant operator, and the exact marginal $(4,0)$. The former indeed tunes the $\mathbb{Z}_{4}$ SSB transition, and the latter moves along a conformal manifold. It is known that the $\mathbb{Z}_{4}$ parafermion CFT can be represented as $\mathrm{U}(1)_{6}/\mathbb{Z}_{2}$, so the conformal manifold is the orbifold line of $c=1$ CFTs.
Via the correspondence, we have a DQCP transition Eq. (101) given by the $\mathbb{Z}_{2}$ orbifold of the $\mathbb{Z}_{4}$ parafermion CFT. It is known that the $\mathbb{Z}_{4}$ parafermion CFT can be represented as $\mathrm{U}(1)_{6}/\mathbb{Z}_{2}$, so the DQCP (multicritical) transition is the $\mathrm{U}(1)_{6}$ CFT. Explicit form of the DQCP partition function is given in Appendix B.
@@ Appendix B $\mathbb{Z}_{N}$ parafermion CFT and the DQCP partition function
$\begin{split}Z_{20}&=Z_{m^{2}}+Z_{em^{2}}+Z_{e^{2}m^{2}}+Z_{e^{3}m^{2}}\\ &=\chi_{4,4}\bar{\chi}_{4,0}+\chi_{1,1}\bar{\chi}_{3,1}+\chi_{3,3}\bar{\chi}_{3,-1}+\chi_{2,2}\bar{\chi}_{2,2}+\chi_{2,0}\bar{\chi}_{2,0}\\ &\quad+\chi_{3,1}\bar{\chi}_{1,1}+\chi_{3,-1}\bar{\chi}_{3,3}+\chi_{4,2}\bar{\chi}_{4,-2}+\chi_{4,0}\bar{\chi}_{4,4}+\chi_{4,-2}\bar{\chi}_{4,2},\end{split}$
$\begin{split}Z_{21}&=Z_{m^{2}}+iZ_{em^{2}}-Z_{e^{2}m^{2}}-iZ_{e^{3}m^{2}}\\ &=-\chi_{4,4}\bar{\chi}_{4,0}-i\chi_{1,1}\bar{\chi}_{3,1}+i\chi_{3,3}\bar{\chi}_{3,-1}+\chi_{2,2}\bar{\chi}_{2,2}-\chi_{2,0}\bar{\chi}_{2,0}\\ &\quad-i\chi_{3,1}\bar{\chi}_{1,1}+i\chi_{3,-1}\bar{\chi}_{3,3}+\chi_{4,2}\bar{\chi}_{4,-2}-\chi_{4,0}\bar{\chi}_{4,4}+\chi_{4,-2}\bar{\chi}_{4,2},\end{split}$
$\begin{split}Z_{22}&=Z_{m^{2}}-Z_{em^{2}}+Z_{e^{2}m^{2}}-Z_{e^{3}m^{2}}\\ &=\chi_{4,4}\bar{\chi}_{4,0}-\chi_{1,1}\bar{\chi}_{3,1}-\chi_{3,3}\bar{\chi}_{3,-1}+\chi_{2,2}\bar{\chi}_{2,2}+\chi_{2,0}\bar{\chi}_{2,0}\\ &\quad-\chi_{3,1}\bar{\chi}_{1,1}-\chi_{3,-1}\bar{\chi}_{3,3}+\chi_{4,2}\bar{\chi}_{4,-2}+\chi_{4,0}\bar{\chi}_{4,4}+\chi_{4,-2}\bar{\chi}_{4,2},\end{split}$
$\begin{split}Z_{23}&=Z_{m^{2}}-iZ_{em^{2}}-Z_{e^{2}m^{2}}+iZ_{e^{3}m^{2}}\\ &=-\chi_{4,4}\bar{\chi}_{4,0}+i\chi_{1,1}\bar{\chi}_{3,1}-i\chi_{3,3}\bar{\chi}_{3,-1}+\chi_{2,2}\bar{\chi}_{2,2}-\chi_{2,0}\bar{\chi}_{2,0}\\ &\quad+i\chi_{3,1}\bar{\chi}_{1,1}-i\chi_{3,-1}\bar{\chi}_{3,3}+\chi_{4,2}\bar{\chi}_{4,-2}-\chi_{4,0}\bar{\chi}_{4,4}+\chi_{4,-2}\bar{\chi}_{4,2}.\end{split}$ (217)